What’s the current yield of a 5.65 percent coupon corporate bond quoted at a price of 97.23?
Answer #1

Annual coupon=1000*5.65%=56.5

Current price=$1000*97.23%=$972.3

Current yield=Annual coupon/Current price

=56.5/972.3

=5.81%(Approx).